# Multistep Inverse Is Not All You Need

### v1

@@ 2.1 Ex-BMDP Model
  • $s\in\mathcal{S}$ and $e\in\mathcal{E}$ are the control-endogenous and control-exogenous latent states. We assume that $\mathcal{S}$ is finite, and typically $|\mathcal{S}|\ll|\mathcal{X}|$. [^2] In our proofs, we also assume that $\mathcal{E}$ and $\mathcal{X}$ are finite, but this is a technical limitation of our theory: see discussion in Appendix B. In any case, the Ex-BMDP formulation is most useful when $|\mathcal{S}|\ll|\mathcal{X}|$.
@@ 2.2 AC-State
Lamb et al. (2022) then proposes to learn the endogenous encoder $\phi_{\theta}$ by learning a multi-step inverse dynamics model. This model is a learned classifier $f(\phi_{\theta}(x_{t}),\phi_{\theta}(x_{t+k});k)$ which takes the encoded endogenous latent states of two observations separated by $k\leq D$ time-steps, as well as $k$, and returns a normalized distribution over actions in $\mathcal{A}$. This classifier is trained to predict the first action $a_{t}$ taken on the trajectory from $x_{t}$ to $x_{t+k}$. The classifier $f$ is trained jointly with the encoder model. Then, in the theoretical treatment, the optimal encoder $\phi_{\theta^{*}}$ is defined as the encoder which allows $f$ to reach the minimum achievable value of this classification loss while also using the fewest number of distinct output states. (In practice, we use a discrete bottleneck and an entropic loss term to minimize the output range, rather than learning multiple models.) Explicitly:
@@ 3.1 Why $D$ Steps is Insufficient for Multi-Step Dynamics
The proofs in Lamb et al. (2022) implicitly assume that the witness distance $W(a,b)$ of any pair of states $a,b\in\mathcal{S}$ is upper-bounded by the diameter $D$ of the transition graph. To briefly sketch the main proof in Lamb et al. (2022), let $W(a,b)=k\leq D$ with witness state $c$. Assume $s_{t}=c$ and $s_{t+k}\in\{a,b\}$, and we wish to accurately predict $a_{t}$ given $s_{t}$ and $s_{t+k}$. Being able to distinguish $a$ and $b$ is guaranteed to help us make this prediction, because the sets of possible values that $a_{t}$ can take will be disjoint depending on if $s_{t+k}=a$ or $s_{t+k}=b$. (Otherwise, the witness distance would be most $k-1$, because $a$ and $b$ could both be reached from $s_{t+1}$, leading to a contradiction.)
However, it is not true in general that $W(a,b)\leq D$. Figure 3-B,C,D gives an example on a simple four-state graph, where $D=3$, but the largest witness distance $W(c,d)=4$. The AC-State loss with $K=D=3$ will learn an encoder that fails to distinguish all of the endogenous latent states. (To analyse this example: we can’t infer the action $a_{t}$ from $s_{t}$ and $s_{t+k}$ if $s_{t}$ is any state other than $a$, because for all other states, the actions L and R have the same effect. From $a$, the states reachable in exactly $k=1$ step are $\{b,d\}$; in $k=2$ steps are $\{c,a\}$; and in $k=3$ steps are $\{a,b,d\}$. In particular, when the classifier $f$ is provided with the value of $k\in\{1,2,3\}$, there is no need for the encoder to be able to distinguish the state $c$ from the state $d$.)
@@ 3.1.1 $D^{\prime}:=2D^{2}+D$ is a Tight Upper-Bound
Furthermore, we show that this bound is tight up to a constant multiplicative factor. We explicitly construct Ex-BMDPs on which AC-State learns an incorrect state encoder for any $K<D^{2}/2+O(D)$, for an infinite sequence of arbitrarily-large values of $D$. (If we use both the multistep-inverse loss and a latent forward-dynamics loss, as in the ACDF algorithm (Equation 3), the particular family of Ex-BMDPs we use to derive this lower-bound no longer minimizes the loss. However, through an alternative construction, we can find Ex-BMDPs where the ACDF loss is minimized by an incorrect encoder if $K<D^{2}/4+O(D)$. Therefore the upper bound $D^{\prime}:=2D^{2}+D$ is still tight for ACDF.)
@@ 5 Related Works
In the area of learning latent representations for reinforcement learning problems, there are numerous ways of defining the “purpose” of the representation – i.e., what information should ideally be included or excluded from the representation. Our work specifically focuses on the Ex-BMDP formulation given in Section 1, which was first proposed by Efroni et al. (2022b), and later studied by Lamb et al. (2022). Note that the setting considered by Efroni et al. (2022b) is time-inhomogeneous: the Ex-BMDP is assumed to progress for a finite number of steps $H$ from a (near) deterministic control-endogenous start state, and the algorithm learns a different state representation for each time-step. Consequently, number of states needed to represent the dynamics is potentially greatly increased, and this setting does not allow for generalization to long sequences, or encoding when the current time step is not known. In contrast, Lamb et al. (2022) and this work consider the infinite-time-horizon Ex-BMDP. Note that because AC-State is the only prior work to our knowledge that considers the infinite-time-horizon finite-$\mathcal{S}$ Ex-BMDP explicitly, we compare to it as our sole baseline in experiments. Other works present multi-step inverse methods learning for timestep-dependent dynamics, including by Mhammedi et al. (2023), which allows for nondeterministic endogenous dynamics, and Efroni et al. (2022a), which considers an explicitly factored state. Multi-step inverse methods have also been used in the continuous-state setting. Mhammedi et al. (2020) considers the special case of linear control-endogenous dynamics, and derives theoretical guarantees in this setting. More recently, Islam et al. (2023) and Koul et al. (2023) have used multi-step inverse methods as an empirical technique for representation learning under continuous latent states. Note that while Koul et al. (2023) does learn a forward dynamics model, the forward-dynamics loss is not used to train the state encoder: it is only used for planning.
@@
Aside from the Ex-BMDP framework and multi-step inverse methods, other methods have been proposed to learn compact relevant state representations. Efroni et al. (2022b) discusses how each of these classes of methods will sometimes include control-exogenous noise into the latent representation, or fail to include control-endogenous information. Techniques such as Deep Bisimulation for Control (Zhang et al., 2020) and DeepMDP (Gelada et al., 2019) ultimately rely on an external reward signal to determine what features are relevant for control, and so may fail to represent controllable aspects of the environment that do not affect the training reward. Misra et al. (2020) demonstrates that one-step inverse dynamics, such as those used empirically by Pathak et al. (2017), are insufficient for learning control-endogenous dynamics. Techniques based on “compressing” states, e.g. through auto-encoders (Hafner et al., 2019) may also include exogenous information.
@@ Appendix B Technical Assumptions
    We assume that, for the exogenous encoder $\phi^{*}_{e}$ corresponding to the above $\phi^{*}$, there are no transient endogenous latent states in $\mathcal{E^{*}}$.
@@ Appendix D Bounds on D’ and K
[…]
@@ Appendix E ACDF returns a correct, minimal Encoder
Consider an Ex-BMDP with states $\mathcal{X}$, transition function $\mathcal{T}$, canonical endogenous states $\mathcal{S}^{*}$ and canonical endogenous states $\mathcal{S}^{*}$, canonical encoders $\phi^{*}$ and $\phi_{e}^{*}$, and canonical emission distribution $\mathcal{Q}^{*}$. If some encoder ϕsuperscriptitalic-$\phi^{\prime}$ exists, such that:
@@
[…]
then ϕsuperscriptitalic-$\phi^{\prime}$ is a valid endogenous latent encoder for the Ex-BMDP.
@@
Given an Ex-BMDP, under the assumptions listed in Section B, there exists a correct minimal-state endogenous state encoder $\phi^{*}$ that will minimize the ACDF loss given in Equation 3 on the Ex-BMDP. Conversely, any encoder ϕsuperscriptitalic-$\phi^{\prime}$ which minimizes the ACDF loss given in Equation 3 on the Ex-BMDP is a correct endogenous state representation for the Ex-BMDP. As a consequence, the encoder ϕsuperscriptitalic-$\phi^{\prime}$ which minimizes Equation 3 with the minimum number of output states is a correct minimal-state endogenous representation.
@@
(2) An encoder ϕsuperscriptitalic-$\phi^{\prime}$ which minimizes both loss terms is a correct endogenous encoder.
@@ H.2 Architecture and Hyperparameters
In general, unless otherwise noted here, we use exactly the architecture and training hyperparameter settings that are default in the released code of Lamb et al. (2022) for gridworld exploration environments.
@@ H.4 Implementation of the forward-dynamics loss
To implement the forward dynamics loss of ACDF (Equation 3), for the dynamics model $g$ we use a LeakyReLU MLP consisting of four layers of sizes $[512+10,1024,1024,N]$ where the input is the discrete-vector output of the VQ-VAE representing the encoded state (a vector of size 512), concatenated with the action label. The output the the discrete index of the VQ-VAE code of the next latent state. (The number of codes is a hyperparameter of AC-State which we vary.) Note that because we treat the next-state as a discrete index, we do not backpropagate into the encoder of $\phi_{\theta}(x_{t+1})$, only into the encoder of $\phi_{\theta}(x_{t})$. This is to mitigate the “moving-target” issue mentioned at the beginning of Appendix F. Also to mitigate this issue, we only use the forward prediction loss to update the encoder $\phi_{\theta}$ every fifth training iteration: at other iterations, we update the parameters of $g$ alone (and separately update $\phi_{\theta}$ with the multistep-inverse loss).
@@ H.6 Results
For both environments and both methods, we ran a hyperparameter sweep over $K$, the number of steps used in multistep-inverse, and $N$, the maximum number of latent states (i.e., the size of the codebook of the VQ-VAE). For each hyperparameter setting, we ran on 10 seeds. We then evaluated on the “best” version of the hyperparameters for each method for 20 additional seeds. Our final results reflect the rate of successful representation learning at the 98% planning success threshold, at training step 30,000, as described above. Note that the final averages are over the 20 additional seeds alone, to avoid multiple-comparison issues.

### v2

@@ 2.1 Ex-BMDP Model
  • $s\in\mathcal{S}$ and $e\in\mathcal{E}$ are referred to as the control-endogenous and control-exogenous latent states, respectively. We assume that $\mathcal{S}$ is finite, and typically $|\mathcal{S}|\ll|\mathcal{X}|$. [^2] In our proofs, we also assume that $\mathcal{E}$ and $\mathcal{X}$ are finite, but this is a technical limitation of our theory: see discussion in Appendix B. In any case, the Ex-BMDP formulation is most useful when $|\mathcal{S}|\ll|\mathcal{X}|$.
@@ 2.2 AC-State
Lamb et al. (2022) then proposes to learn the endogenous encoder $\phi_{\theta}$ by learning a multi-step inverse dynamics model. This model is a learned classifier $f(\phi_{\theta}(x_{t}),\phi_{\theta}(x_{t+k});k)$ which takes the encoded endogenous latent states of two observations separated by $k\leq D$ time-steps, as well as $k$, and returns a normalized distribution over actions in $\mathcal{A}$. This classifier is trained to predict the first action $a_{t}$ taken on the trajectory from $x_{t}$ to $x_{t+k}$. The classifier $f$ is trained jointly with the encoder model. Then, in the theoretical treatment, the optimal encoder $\phi_{\theta^{*}}$ is defined as the encoder which allows $f$ to reach the minimum achievable value of this classification loss while also using the fewest number of distinct output states. (In practice, Lamb et al. (2022) uses a discrete information bottleneck and an associated loss term to minimize the output range, rather than learning multiple models.) Explicitly:
@@ 3.1 Why $D$ Steps is Insufficient for Multi-Step Dynamics
The proofs in Lamb et al. (2022) implicitly assume that the witness distance $W(a,b)$ of any pair of states $a,b\in\mathcal{S}$ is upper-bounded by the diameter $D$ of the transition graph. To briefly sketch the main proof in Lamb et al. (2022), let $W(a,b)=k\leq D$ with witness state $c$. Assume $s_{t}=c$ and $s_{t+k}\in\{a,b\}$, and we wish to accurately predict $a_{t}$ given $x_{t}$ and $x_{t+k}$ (in order to minimize the $k$-step inverse loss, which is part of our loss function because $k\leq D$, by assumption). Being able to determine from $x_{t+k}$ whether $s_{t+k}=a$ or $s_{t+k}=b$ is guaranteed to help us make this prediction, because the sets of possible values that $a_{t}$ can take will be disjoint depending on if $s_{t+k}=a$ or $s_{t+k}=b$. (Otherwise, the witness distance would be most $k-1$, because $a$ and $b$ could both be reached from $s_{t+1}$, leading to a contradiction.)
However, it is not true in general that $W(a,b)\leq D$. Figure 3-B,C,D gives an example on a simple four-state graph, where $D=3$, but the largest witness distance $W(c,d)=4$. The AC-State loss with $K=D=3$ will learn an encoder that fails to distinguish all of the endogenous latent states. To analyse this example: we can’t infer the action $a_{t}$ from $s_{t}$ and $s_{t+k}$ if $s_{t}$ is any state other than $a$, because for all other states, the actions L and R have the same effect. From $a$, the states reachable in exactly $k=1$ step are $\{b,d\}$; in $k=2$ steps are $\{c,a\}$; and in $k=3$ steps are $\{a,b,d\}$. In particular, when the classifier $f$ is provided with the value of $k\in\{1,2,3\}$, there is no need for the encoder to be able to distinguish the state $c$ from the state $d$.
@@ 3.1.1 $D^{\prime}:=2D^{2}+D$ is a Tight Upper-Bound
Furthermore, we show that this bound is tight up to a constant multiplicative factor. We explicitly construct Ex-BMDPs on which AC-State learns an incorrect state encoder for any $K<D^{2}/2+O(D)$, for an infinite sequence of arbitrarily-large values of $D$. If we use both the multistep-inverse loss and a latent forward-dynamics loss, as in the ACDF algorithm (Equation 3), the particular family of Ex-BMDPs we use to derive this lower-bound no longer minimizes the loss. However, through an alternative construction, we can find Ex-BMDPs where the ACDF loss is minimized by an incorrect encoder if $K<D^{2}/4+O(D)$. Therefore the upper bound $D^{\prime}:=2D^{2}+D$ is still tight for ACDF.
@@ 5 Related Works
In the area of learning latent representations for reinforcement learning problems, there are numerous ways of defining the “purpose” of the representation – i.e., what information should ideally be included or excluded from the representation. Our work specifically focuses on the Ex-BMDP formulation given in Section 1, which was first proposed by Efroni et al. (2022b), and later studied by Lamb et al. (2022). Note that the setting considered by Efroni et al. (2022b) is time-inhomogeneous: the Ex-BMDP is assumed to progress for a finite number of steps $H$ from a (near) deterministic control-endogenous start state, and the algorithm learns a different state representation for each time-step. Consequently, number of states needed to represent the dynamics is potentially greatly increased, and this setting does not allow for generalization to long sequences, or encoding when the current time step is not known. In contrast, Lamb et al. (2022) and this work consider the infinite-time-horizon Ex-BMDP. Note that because AC-State is the only prior work to our knowledge that considers the infinite-time-horizon finite-$\mathcal{S}$ Ex-BMDP explicitly, we compare to it as our sole baseline in experiments. (While Wang et al. (2022) considers a similar formulation, with continuous latent states $\mathcal{S}$, the approach proposed in Wang et al. (2022) requires learning a generative model of the complete system state $\mathcal{X}$, explicitly modeling even the exogenous state $\mathcal{E}$ of the system. Kooi et al. (2023) also requires explicitly modeling both the “controllable” and “uncontrollable” latent dynamics, and also uses a somewhat different definition for how these factors relate than in the Ex-BMDP framework.)
Other works present multi-step inverse methods learning for timestep-dependent dynamics, including by Mhammedi et al. (2023), which allows for nondeterministic endogenous dynamics, and Efroni et al. (2022a), which considers an explicitly factored state. Multi-step inverse methods have also been used in the continuous-state setting. Mhammedi et al. (2020) considers the special case of linear control-endogenous dynamics, and derives theoretical guarantees in this setting. More recently, Islam et al. (2023) and Koul et al. (2023) have used multi-step inverse methods as an empirical technique for representation learning under continuous latent states. Note that while Koul et al. (2023) does learn a forward dynamics model, the forward-dynamics loss is not used to train the state encoder: it is only used for planning.
@@
Aside from the Ex-BMDP framework and multi-step inverse methods, other methods have been proposed to learn compact relevant state representations. Efroni et al. (2022b) discusses how several of these classes of methods will sometimes include control-exogenous noise into the latent representation, or fail to include control-endogenous information. Techniques such as Deep Bisimulation for Control (Zhang et al., 2020) and DeepMDP (Gelada et al., 2019) ultimately rely on an external reward signal to determine what features are relevant for control, and so may fail to represent controllable aspects of the environment that do not affect the training reward. Misra et al. (2020) demonstrates that one-step inverse dynamics, such as those used empirically by Pathak et al. (2017), are insufficient for learning control-endogenous dynamics. Techniques based on “compressing” states, e.g. through auto-encoders (Hafner et al., 2019) may also include exogenous information.
@@ Appendix B Technical Assumptions
    We assume that, for the exogenous encoder $\phi^{*}_{e}$ corresponding to the above $\phi^{*}$, there are no transient exogenous latent states in $\mathcal{E^{*}}$.
@@ Appendix D Bounds on D’ and K
Consider any set of positive integers =$\mathcal{B}=\{b_{1},...,b_{k}\}$ such that $\text{gcd}(\mathcal{B})=1$, where $k=|\mathcal{B}|$, $b_{1}$ is the smallest integer, and $b_{k}$ the largest. Then any integer $n\geq(b_{1}-1)(b_{k}-1)$ can be represented as a sum in the form $n=\sum_{i=1}^{k}a_{i}b_{i}$, where $a_{1},...,a_{k}$ are non-negative integers.
@@ Appendix E ACDF returns a correct, minimal Encoder
Consider an Ex-BMDP with states $\mathcal{X}$, transition function $\mathcal{T}$, canonical endogenous states $\mathcal{S}^{*}$ and canonical endogenous states $\mathcal{S}^{*}$, canonical encoders $\phi^{*}$ and $\phi_{e}^{*}$, and canonical emission distribution $\mathcal{Q}^{*}$. If some encoder $\phi^{\prime}$ exists, such that:
@@
    The dynamics on the encoded latent states produced by $\phi^{\prime}(x)$ are deterministic. That is, if we fix any $s\in\text{Range}(\phi^{\prime})$ and $a\in\mathcal{A}$, then $\forall x\in\{x|\phi^{\prime}(x)=s\},$ let $x^{\prime}\sim\mathcal{T}(x,a)$, then $\phi^{\prime}(x^{\prime})$ takes a single deterministic value, a function only of $s$ and $a$. Additionally, we assume that the range of $\phi^{\prime}$ is discrete and finite.
then $\phi^{\prime}$ is a valid endogenous latent encoder for the Ex-BMDP.
@@
Given an Ex-BMDP, under the assumptions listed in Section B, there exists a correct minimal-state endogenous state encoder $\phi^{*}$ that will minimize the ACDF loss given in Equation 3 on the Ex-BMDP. Conversely, any encoder $\phi^{\prime}$ which minimizes the ACDF loss given in Equation 3 on the Ex-BMDP is a correct endogenous state representation for the Ex-BMDP. As a consequence, the encoder $\phi^{\prime}$ which minimizes Equation 3 with the minimum number of output states is a correct minimal-state endogenous representation.
@@
(2) An encoder $\phi^{\prime}$ which minimizes both loss terms is a correct endogenous encoder.
@@ H.2 Architecture and Hyperparameters
In general, unless otherwise noted here, we use exactly the architecture and training hyperparameter settings that are default in the released code of Lamb et al. (2022) for gridworld exploration environments. [^8] In particular, we used exactly the network architecture for the encoder provided by Lamb et al. (2022)’s released code, for both AC-State and ACDF. We have observed that this architecture design appears to be specifically well-suited to easily learn to filter out the exogenous “distractor” mazes in these multi-maze environments, and may not be more generally applicable. However because we use the same architecture for both ACDF and AC-State, our direct comparison of these methods is valid.
@@ H.4 Implementation of the forward-dynamics loss
To implement the forward dynamics loss of ACDF (Equation 3), for the dynamics model $g$ we use a LeakyReLU MLP consisting of four layers of sizes $[512+10,1024,1024,N]$ where the input is the discretized vector output of the vector-quantization output layer of the state encoder, representing the encoded state (a vector of size 512), concatenated with the action label. The output is the discrete index of the vector-quantized latent representation code of the next latent state. (The number of codes is a hyperparameter of AC-State which we vary.) Note that because we treat the next-state as a discrete index, we do not backpropagate into the encoder of $\phi_{\theta}(x_{t+1})$, only into the encoder of $\phi_{\theta}(x_{t})$. This is to mitigate the “moving-target” issue mentioned at the beginning of Appendix F. Also to mitigate this issue, we only use the forward prediction loss to update the encoder $\phi_{\theta}$ every fifth training iteration: at other iterations, we update the parameters of $g$ alone (and separately update $\phi_{\theta}$ with the multistep-inverse loss).
@@ H.6 Results
For both environments and both methods, we ran a hyperparameter sweep over $K$, the number of steps used in multistep-inverse, and $N$, the maximum number of latent states (i.e., the size of the codebook of the final vector-quantization layer of the encoder). For each hyperparameter setting, we ran on 10 seeds. We then evaluated on the “best” version of the hyperparameters for each method for 20 additional seeds. Our final results reflect the rate of successful representation learning at the 98% planning success threshold, at training step 30,000, as described above. Note that the final averages are over the 20 additional seeds alone, to avoid multiple-comparison issues.